My problem involves solving quadratic equations using the zero product property.
I couldn't figure out how to make this one work. It is: m^3+5m^2-9m-45=0. Any help would be greatly appreciated.

Answer by scott8148(6628)

m^3+5m^2-9m-45=0 is not a quadratic , it is a cubic

grouping ___ (m^3 + 5m^2) - (9m + 45) = 0

factoring ___ m^2(m + 5) - 9(m + 5) = 0

rearranging ___ (m^2 - 9)(m + 5) = 0

factoring (again) ___ (m + 3)(m - 3)(m + 5) = 0